生体機能工学 2 単位 (選択)

Biofunctional Engineering  松木均 · 教授/ 環境創生工学専攻生命テクノサイエンスコース生物機能工学講座

【授業目的】細胞の構造形成や機能発現に大きな役割を果たしているのは生体膜 である．脂質膜が有する特徴的な物性を中心に講述し，生体膜の構造と機能 について解説する． 【授業概要】生体分子の自己会合に関する基本的な事項 (水の構造と疎水効果，両 親媒性分子の分子集合体特性) を述べた後に，脂質が形成する分子集合体が示 す様々な構造的変化 (相転移，多形現象，非二重膜構造，ドメイン形成) につ いて講述する．さらに輸送現象，薬物作用機構など生体膜を場として発現す る重要な機能について述べる．受講学生によっては，ポートフォリオ形式を 併用する． 【授業形式】講義形式とポートフォーリオ形式の併用 【キーワード】生体膜，脂質，分子集合体，膜構造，膜機能 【先行科目】『生体熱力学』 (1.0)，『生物物理化学特論』 (1.0) 【関連科目】 [関連科目] 【履修要件】学部および大学院教育における物理化学および生物物理化学を理解 していること． 【履修上の注意】 [注意] 【到達目標】 1. 生体膜構成成分の脂質が形成する分子集合体構造とその性質を理解する． 2. 生体膜により発現する機能 (物質輸送，情報伝達，薬物作用) を理解する． 【授業計画】 1. 生体膜の構造 (1) 水の構造と疎水性相互作用 2. 生体膜の構造 (2) 両親媒性物質の自己会合 1:単分子膜 3. 生体膜の構造 (3) 両親媒性物質の自己会合 2:ミセル 4. 生体膜の構造 (4) 脂質膜の二重膜構造と相転移 5. 生体膜の構造 (5) 脂質膜の安定性と多形現象 6. 生体膜の構造 (6) 脂質膜の非二重膜構造 7. 生体膜の構造 (7) 混合脂質膜とドメイン形成 8. 生体膜の構造 (8) 脂質膜の流動化とダイナミクス 9. 生体膜の構造 (9) 表面電位と膜電位 10. 生体膜の機能 (1) 脂質膜の分子認識 11. 生体膜の機能 (2) イオン輸送 (受動および能動輸送) 12. 生体膜の機能 (3) イオンチャンネルと神経伝導 13. 生体膜の機能 (4) 薬物作用機構 (特異的結合)

14. 生体膜の機能 (5) 薬物作用機構 (非特異的結合)

15. 生体膜の機能 (6) 細胞機能とシグナル伝達

16. 生体膜と構造と機能に関するレポート作成【成績評価基準】課題レポートにより評価する． (100%) 【教科書】授業中に紹介する. 【参考書】授業中に紹介する．【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216728 【対象学生】開講コース学生のみ履修可能【連絡先】

⇒ 松木 (化生棟 607, 088-656-7513, matsuki@bio.tokushima-u.ac.jp) ���� (オ

フィスアワー: 金曜日 16:20-17:50) 【備考】到達目標 1 は授業計画 1-9 に，到達目標 2 は授業計画 10-15 に関係する．

Biofunctional Engineering 2 units (selection)

Hitoshi Matsuki · Professor / Biological Functions, Biological Science and Technology, Earth and Life Environmental Engineering

Target⟩ Biological membranes play an important roll in structure formation and function revelation of cells. Characteristic properties owned by lipid membranes are mainly described and the structure and function of biological membranes are reviewed.

Outline⟩ This lecture describes fundamental matters concerning self-association of biomolecules such as structure of water and hydrophobic effect, properties for molecular aggregates of amphiphilic molecules, after that various structure changes for molecular aggregates formed by lipids such as phase transitions, polymorphism, non-bilayer structure and domain formation are explained. Further, significant functions revealed in biological membranes like transport phenomena and mechanisms of drug action and so on are lectured. Portfoilo lessons are also adapted depending on students.

Style⟩ Lecture in combination with Portfolio �������⟩ biological membrane, lipid, molecular aggregate, membrane structure,

membrane function Fundamental Lecture⟩ “Biochemical Thermodynamics”(1.0), “Advanced Bio-

physical Chemistry”(1.0) Relational Lecture⟩ [関連科目] Requirement⟩ Students are required to have a good understanding of undergraduate-

and postgraduate-level physical chemistry and biophysical chemistry and related subjects.

Notice⟩ [注意]

Goal⟩

1. To understand the structures and nature of molecular aggregates formed by lipids, which are components of biological membranes.

2. To understand the functions revealed by biological membranes such as

material transport, signal transduction and drug action.

Schedule⟩

1. Structures of biological membranes (1) structure of water and hydrophobic interaction

2. Structures of biological membranes (2) self-association of amphiphilic molecules 1: monolayers 3. Structures of biological membranes (3) self-association of amphiphilic

molecules 2: micelles

4. Structures of biological membranes (4) bilayer structure and phase transitions of lipid membranes

5. Structures of biological membranes (5) stability and polymorphism of lipid membranes

6. Structures of biological membranes (6) non- bilayer structure of lipid membranes

7. Structures of biological membranes (7) mixed lipid membranes and domain formation

8. Structures of biological membranes (8) fluidization and dynamics of lipid membranes

9. Structures of biological membranes (9) surface and membrane potentials

10. Functions of biological membranes (1) molecular recognition of lipid membranes

11. Functions of biological membranes (2) ion transport (passive and active transport)

12. Functions of biological membranes (3) ion channel and nerve conduction 13. Functions of biological membranes (4) mechanisms of drug action (specific

binding)

14. Functions of biological membranes (5) mechanisms of drug action (non-specific binding

15. Functions of biological membranes (6) cell functions and signal transduction 16. Report preparation for structures and functions of biological membranes Evaluation Criteria⟩ Assignments count 100%. Textbook⟩ To be introduced in the class. Reference⟩ To be introduced in the class.

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216728

Student⟩ Able to be taken by only specified class(es)

Contact⟩

⇒ Matsuki (G607, +81-88-656-7513, matsuki@bio.tokushima-u.ac.jp) ����

(Office Hour: Friday 16:20-17:50)

Note⟩ Goal 1 is related to schedules 1-9 and goal 2 is related to schedules 10-15.